5. Solve the equation 7 +2+y = 2+32) (do tiot use the Laplace transform.) 6. Use the definition of Laplace transform to findthis is from differential equations. if someone can do the whole sheet except for question 10 if you could do two to of them, I would greatly appreciate it. thank you in advance.
Answer #1

Given yapzyty= ē^ca+3n) Cousider the Homojeneoms problem y+zy+y=o -. Auni liary equation is +2+120 soluing weget Y= ų = qit ostei fr+)= ,it t?, 3 తైి fet)dt fe.. %3D %3D - --*) +(* +) + (0-e ) (y1) endn+ 3yCe+) dy=0 Comparting corith po cu+ Ndy